# Can You Trust LLM Judgments? Reliability of LLM-as-a-Judge

### v1

@@ 5 Experimental Results
Table 1 presents a comprehensive overview of judgment reliability for temperature variations across various benchmarks and LLMs. While one might expect that decreasing temperature would lead to a corresponding increase in reliability, as observed with a temperature of 0 (resulting in identical outputs), the reality is more intricate.
Our findings reveal a complex interplay between temperature and reliability. While some cases, such as Gemma-1.1-7b-it judgments of SQuAD responses (Table 1(c)), exhibit a clear decrease in reliability with increasing temperature, others, like Meta-Llama-3-8B-Instruct judgments of BBH prompts (Table 1(b)), show the opposite trend. This unexpected behavior highlights the nuanced relationship between temperature and LLM output quality. While higher temperatures can indeed lead to more novel and diverse outputs, they can also introduce variability and potential degradation in reliability. The ability to maintain reliability while increasing temperature is a promising avenue for future research, as it opens up opportunities to explore methods that generate novel, reliable, and high-quality outputs.
@@
While the reliability across temperatures shows promise, the reliability values themselves in Table 1 are a cause of concern, highlighting significant inconsistencies in judgments across replications. It’s important to note that the calculated reliability values are an optimistic estimate of the true reliability, as non-response judgments were treated as belonging to the same category. This assumption likely overestimates the reliability, particularly for models with high rates of non-response. Adhering to the widely accepted rule of thumb in Table 2 for interpreting reliability values in practice, we assess the reliability of our judgments (Gliem et al., 2003). Starling-LM-7B-beta and Meta-Llama-3-8B-Instruct exhibit questionable, at best, reliability across all benchmarks, with SQuAD and MT-Bench results consistently falling below the acceptable range (Tables 1(a) and 1(b), respectively). This low reliability raises concerns about the generalizability of their judgments. Gemma-1.1-7b-it fares better, with BBH judgments falling within the acceptable range or higher across all temperatures. SQuAD results are predominantly within the acceptable range, and no MT-Bench results fall within the unacceptable range (Table 1(c)). This makes Gemma-1.1-7b-it our most reliable judge, followed by Starling-LM-7B-beta and then Meta-Llama-3-8B-Instruct. Interestingly, this reliability ranking is the inverse of the models’ performance on the Chatbot Arena. This suggests a potential trade-off between model performance and judgment reliability, where models prioritizing performance may sacrifice reliability in their assessments.
@@
To accurately assess benchmark performance, we must consider the inherent quality of response options [A]-[E]. As shown in Table 3, SQuAD questions exhibit significantly and consistently higher response accuracy compared to BBH questions. This distinction is pivotal. Lower response accuracy makes it easier to identify a clear best answer, as the difference in quality between options is more pronounced. In contrast, higher accuracy often results in multiple correct options, demonstrating judges’ difficulty in discerning the best response based solely on explanations. Consequently, the lower reliability observed for SQuAD across models and temperatures in Table 1 likely reflects this inherent difficulty in judging among highly accurate responses. This result is important to take into consideration in the employment of LLM-as-a-judge, as it points to potential struggle of LLM judges to differentiate between subtly nuanced correct responses.
Table 1 also shows that reliability is consistently worse for MT-Bench than for BBH and often worse for MT-Bench compared to SQuAD. Accuracies for MT-Bench are not provided in Table 3 as MT-Bench questions are open-ended. The subjectivity inherent in evaluating open-ended responses makes it more challenging to achieve consistent judgments, particularly for LLMs which may lack the nuanced understanding required to assess such responses.
@@ 6 Application
We explore the influence of LLM reliability on practical applications by focusing on how LLMs are used as judges in the Head-to-Tail benchmark in Sun et al. (2023). This benchmark aims to assess the knowledge level of LLMs by creating a set of questions about entities with varying levels of popularity ("head" - least popular, "torso" - moderately popular, "tail" - most popular). Focusing on the "head" questions, we replicate the benchmark for the Academic domain using the Vicuna-7B model.
Similar to Sun et al. (2023), we then use another LLM (Starling-LM-7B-beta, Gemma-1.1-7b-it, or Meta-Llama-3-8B-Instruct) to judge the accuracy of the benchmark’s answers, a metric termed $A_{LM}$. As we will show, this highlights reliability issues that ensue when relying on LLMs for evaluation.
6.1 Accuracy
To ensure our work is an accurate representation of Sun et al. (2023), we replicated the $A_{LM}$ results reported for the "head" dataset using the Vicuna-7B model. The $A_{LM}$ metric is calculated by the proportion of answers judged to be "correct" by the LLM judgment model. Therefore, the expected judgment responses on the accuracy of the Vicuna-7B responses are "correct" or "incorrect", with a judgment non-response considered to fall within the "incorrect" category as well. As shown in Table 3, a single instantiation of our judgment models (Starling-LM-7B-beta, Gemma-1.1-7b-it, and Meta-Llama-3-8B-Instruct) can closely replicate the original $A_{LM}$ values, demonstrating consistency across Sun et al. (2023) and our implementation.
$\mathbf{A_{LM}}$ Vicuna-7B
@@ 6.1 Accuracy
Starling-LM-7B-beta 0.0387
Meta-Llama-3-8B-Instruct 0.0397
Gemma-1.1-7b-it 0.0407
Table 4: The head $\mathbf{A_{LM}}$ as calculated by our judgment models Starling-LM-7B-beta, Gemma-1.1-7b-it, and Meta-Llama-3-8B-Instruct in a single instantiation compared to the head $\mathbf{A_{LM}}$ in Sun et al., 2023 for the Vicuna-7B model.
Figure 3 illustrates the variability in $A_{LM}$ across multiple instantiations of LLM judgments. Variation is similar across judgment models, signaling that there is no LLM judge that outperforms the other judges for this application. However, the figure reveals a notable level of variation across instantiations for all models. This variability can be attributed to inherent randomness in the LLM generation process, similar to the concept of error bars in traditional statistical analysis. In this context, the variability across instantiations results in up to a $1\%$ increase or $1\%$ decrease in the $A_{LM}$ metric, impacting the overall understanding of accuracy achieved by a model responding to the Head-to-Tail benchmark questions.
Figure 3: The variation in $A_{LM}$ of Vicuna-7B responses to Head questions across instantiations of judgment models.
6.2 Reliability
@@
Our reliability framework sheds essential insights into the variability observed in LLM judgments. Reliability of LLM judgments across replications in this context, as seen in Table 5, is quite high. It is unsurprising that the Head-to-Tail reliability is much higher than that of our experiments given that the LLM judgments in the Head-to-Tail paradigm only have two options (’correct’ or ’incorrect’) and are provided with a ground truth in the prompt. Thus, this employment of LLM-as-a-judge is much more simplistic than our experiments and would be expected to show a higher reliability. Despite high reliability scores, Figure 3 shows that $A_{LM}$ varies across replications by ±$\pm 1\%$ as previously discussed. This underscores the potential for substantial error in downstream research that utilizes LLM-based evaluations, even when reliability is high across LLM-as-a-judge-replications. To illustrate the practical implications of this variability, consider a scenario where an LLM is used to rank different models based on their performance on a given task. A $2\%$ variation in the LLM’s judgment could lead to significant shifts in the ranking, potentially misrepresenting the true performance of the models. In the context of high-stakes applications such as ranking models to be used for medical diagnosis, even a small variation in judgment could have serious consequences. For example, an LLM-based diagnostic tool that incorrectly classifies a patient’s condition could lead to incorrect treatment and adverse health outcomes. As the stakes of LLM judgments increase, understanding and mitigating this variability becomes paramount.
7 Discussion
Current utilization of LLMs as judges often relies on single outputs as a clear delineation of an LLM’s choice in judgment, leaning upon deterministic settings (Ouyang et al., 2023; Wei et al., 2024). Using deterministic settings, however, leads to fixed randomness which masks inherent variability and fails to adequately encompass the true nature of the judgment responses, presenting a facade of reliability where one does not actually exist.
Akin to confidence intervals or standard errors typically reported in research, our proposed reliability metric for the LLM-as-a-judge paradigm provides a method for measuring uncertainty of the resulting LLM judgments and should be reported alongside LLM-as-a-judge results. To promote user understanding of model outputs, it’s crucial to maintain skepticism about the LLM’s inner workings and to be mindful of potential uncertainties, just as one would with any other model. Reliability bridges this gap. Researchers can gain a more accurate understanding of conclusions derived from LLM-as-a-judge results by reporting reliability alongside these LLM-based results.
Even for simple tasks with high reliability, LLMs can exhibit significant variability in their outputs, as illustrated in Figure 3. This variability, further emphasized by the low reliability values observed in more complex judgment tasks (Table 1), underscores the importance of acknowledging the inherent uncertainty in LLM judgments. Given the variability seen in downstream applications (Figure 3) in the presence of high reliability values (Table 5), variability across replications is particularly concerning for complex judgment setups, beyond basic correctness assessments, that are likely to be used in practice. This variability highlights the importance of informing future research with reliability. As LLMs are increasingly employed for more complex and subjective tasks, such as making critical medical decisions or assessing another LLMs’ handling of social media moderation, understanding and mitigating uncertainty becomes even more crucial. By incorporating reliability into LLM-based research, we can foster a more cautious and responsible approach to LLM usage, mitigating risks and ensuring responsible LLM employment. This includes encouraging critical evaluation of LLM judgment results and promoting transparency and accountability in LLM-based research. This will help ensure that LLMs are used effectively and ethically, maximizing their potential benefits while minimizing their risks.
8 Future Directions
Clearly, reliability is a critical factor in the successful implementation of the LLM-as-a-judge paradigm. Our work provides a foundational framework for assessing LLM evaluation reliability. To ensure responsible LLM usage, future research should develop domain-specific guidelines for interpreting reliability metrics, as different fields have varying tolerance levels for uncertainty.
Further, the impact of prompt engineering on LLM reliability has been widely discussed (Shah, 2024; Wang et al., 2024). As LLMs are susceptible to adversarial prompting (Zhu et al., 2023b), it’s crucial to assess their reliability across multiple instantiations under various prompts. This knowledge will inform improvements in prompt engineering and LLM model robustness, driving advancements in LLM development.
@@ 9 Limitations
While this work represents a significant step forward in understanding and assessing the reliability of LLMs, it is important to acknowledge its limitations. Our work is currently focused on single and multi-turn benchmarks. These benchmarks were selected to provide a robust and generalizable understanding of the impact of reliability under varying and widely used contexts, including both structured and open-ended tasks. Future research can extend our approach to additional datasets and benchmarks. Further, as discussed in Section 8, developing domain-specific practical guidelines for interpreting and applying reliability metrics is an important area for future work given that different domains have varying tolerances for variability. For instance, fields like advertising may be more tolerant of variability, while fields like medicine require much higher levels of certainty. By providing domain-specific guidance, we can ensure that LLMs are used responsibly and effectively in various applications.

### v2

@@ 6 Experimental Results
Our findings reveal several key insights. First, we observed that the ideal temperature for reliable LLM judgment is not universal; it varies depending on the specific LLM and the task, as revealed in Table 2. For example, some models (like Gemma-1.1-7b-it on SQuAD) exhibit decreased reliability at higher temperatures, while others (like Meta-Llama-3-8B-Instruct on BBH) show the opposite trend. Consequently, careful temperature tuning is crucial for each LLM and task to balance output novelty and judgment reliability. The ability to maintain reliability while increasing temperature, however, is a promising avenue for future research, as it provides opportunities to explore methods that generate novel, reliable, and high-quality outputs.
@@
Furthermore, the results in Table 2 highlight that LLM judgment reliability is a major concern. It’s important to note that the calculated reliability values are likely optimistic estimates of the true reliability, as judgment outputs that did not select a response were treated as belonging to the same category (’Non-response’). Despite this, the overall reliability scores are troubling. Adhering to the widely accepted reliability rule of thumb (Table 3), Starling-LM-7B-beta and Meta-Llama-3-8B-Instruct exhibited questionable reliability across benchmarks, with SQuAD and MT-Bench results consistently falling below the acceptable range (Tables 2(a) and 2(b)). This raises serious questions about the trustworthiness of their judgments.
Moreover, our results suggest a potential performance-reliability trade-off. The ranking of LLM judgment reliability (Gemma-1.1-7b-it > Starling-LM-7B-beta > Meta-Llama-3-8B-Instruct) inverts their Chatbot Arena performance ranking. This suggests that models optimized for performance may sacrifice reliability in their evaluations.
@@
Finally, we find that benchmark quality impacts judgment reliability. SQuAD, despite its higher overall answer accuracy (Table 4), paradoxically exhibits lower judgment reliability than BBH (Table 2). We attribute this to the difficulty in distinguishing the best answer among multiple highly accurate responses in SQuAD. This highlights a critical weakness of current LLM judges: struggling with nuanced distinctions between correct answers.
@@ 7 Application
We explore the influence of LLM reliability on practical applications, focusing on how LLMs function as judges within the Head-to-Tail benchmark (Sun et al., 2023). This benchmark assesses LLM knowledge by posing questions about entities of varying popularity ("head," "torso," "tail"). We replicated the benchmark’s "head" question set for the Academic domain using Vicuna-7B, employing Starling-LM-7B-beta, Gemma-1.1-7b-it, and Meta-Llama-3-8B-Instruct as judges. Our goal is to demonstrate the strong connection between high reliability and consistent evaluation.
We replicated the $A_{LM}$ metric (accuracy judged by an LLM) reported in Sun et al. (2023) using Vicuna-7B. Table 5 shows close agreement between our replicated $A_{LM}$ values and the original results, confirming the reliability and reproducibility of LLM-based evaluations in this structured task. The consistent performance across different LLM judges further supports this.
Critically, our analysis reveals a strong link between high reliability and stable evaluation. Table 6 shows remarkably high reliability scores for our LLM judges across Head-to-Tail replications. Similarly, it is clear from the standard errors in the $A_{LM}$ estimates in Table 5 that minimal variability exists across replications. This demonstrates that highly reliable LLM judges produce consistent and trustworthy evaluations, further supporting the validity of reliability.
Judge $\mathbf{A_{LM}}$ Estimate
@@
(N/A)
Starling-LM-7B-beta 0.0396
Meta-Llama-3-8B-Instruct 0.0395
Gemma-1.1-7b-it 0.0395
Table 5: Average $A_{LM}$ of Vicuna-7B responses to Head questions across replications of judgment models as compared to the results reported by Sun et al., 2024. Corresponding standard errors are provided beneath respective estimates for our judgment models.
@@
It is unsurprising that the Head-to-Tail reliability is much higher than that of our experiments given that the LLM judgments in the Head-to-Tail paradigm only have two options (’correct’ or ’incorrect’) and are provided with a ground truth in the prompt. Thus, this employment of LLM-as-a-judge is much more simplistic than our experiments and would be expected to show a higher reliability. While Head-to-Tail uses a simplified paradigm, it offers valuable insight into the reliability-consistency relationship. The observed high reliability, coupled with minimal variation in $A_{LM}$, strongly suggests that with careful design and clear criteria, LLMs can be dependable judges. Future research should extend these findings to more complex scenarios, developing methods to structure complex judgment tasks to leverage LLM reliability and enhance trustworthiness in diverse applications. Our findings reinforce reliability’s importance in trustworthy LLM evaluation and highlight the potential for reliable LLM judgment in well-defined tasks.
8 Conclusions
Current LLM-as-a-judge methods, relying on single outputs, mask inherent judgment variability, creating a false sense of reliability (Ouyang et al., 2023; Wei et al., 2024). This "fixed randomness" (Section 1), poses risks, especially in high-stakes applications like medical AI evaluation. We address this by introducing a novel framework for evaluating LLM-as-a-judge reliability, focusing on internal consistency across replicated evaluations with varying random seeds — a crucial, previously overlooked aspect.
Our results validate our framework’s importance. LLM judgment reliability is a significant concern, with several models exhibiting questionable reliability across benchmarks like SQuAD and MT-Bench. SQuAD’s low reliability, despite high accuracy, reveals LLMs’ difficulty with nuanced distinctions. Our research also suggests a performance-reliability trade-off, which our framework helps navigate. Finally, Head-to-Tail results further demonstrate our framework’s value, as high reliability scores and minimal variability highlight the link between reliability and consistent evaluation.
This work provides a robust framework for assessing LLM-as-a-judge reliability, addressing a critical gap and developing a crucial tool for building trust in LLM-powered systems. As LLMs are increasingly used in complex applications, reliably assessing their judgment is paramount. Our framework empowers informed decisions about LLM deployment, promoting responsible usage. Prioritizing reliability is key to unlocking LLMs’ potential while ensuring their ethical use.
@@ 9 Limitations
While this work represents a significant step forward in understanding and assessing the reliability of LLMs, it is important to acknowledge its limitations. Our work is currently focused on single and multi-turn benchmarks. These benchmarks were selected to provide a robust and generalizable understanding of the impact of reliability under varying and widely used contexts, including both structured and open-ended tasks. Future research can extend our approach to additional datasets and benchmarks. Further, developing domain-specific practical guidelines for interpreting and applying reliability metrics is an important area for future work given that different domains have varying tolerances for variability. For instance, fields like advertising may be more tolerant of variability, while fields like medicine require much higher levels of certainty. By providing domain-specific guidance, we can ensure that LLMs are used responsibly and effectively in various applications.